# Data Augmentation for Supervised Graph Outlier Detection with Latent Diffusion Models

### v2

@@ III-C Latent Diffusion Model
Forward Diffusion Process. We construct a forward diffusion process $\{\mathbf{Z}(s)\}_{s=0}^{S}$, where $s\in[0,S]$ is a continuous diffusion step. In the diffusion process, $\mathbf{Z}(0)=\mathbf{Z}$ is the embedding from Variational Encoder, while $\mathbf{Z}(S)\sim\mathcal{N}(\bm{0},\bm{I})$ is sampled from the prior distribution. According to [23], the forward diffusion process can be written in stochastic differential equation (SDE) as:
@@
where $\bm{\omega}_{t}$ is the standard Wiener process (i.e., Brownian motion). $\bm{f}(\cdot,s)$ and $g(\cdot)$ are the drift coefficient and the diffusion coefficient, respectively. The selection of $\bm{f}(\cdot,s)$ and $g(\cdot)$ vary between different diffusion models. $f(\cdot)$ is usually of the form $\bm{f}(\mathbf{Z},s)=f(s)\mathbf{Z}$. Thus, Eq. 15 can be written as:
@@
Let $\mathbf{Z}$ be a function of diffusion step $s$, i.e., $\mathbf{Z}^{s}=\mathbf{Z}(s)$. The diffusion kernel of Eq. 16 can be represented in the conditional distribution of $\mathbf{Z}^{s}$ given $\mathbf{Z}^{0}$:
@@
$a(s)=\exp{\int_{0}^{s}f(\xi)\rm d\xi},$ (18)
$\sigma(s)=\sqrt{\int_{0}^{s}g^{2}(\xi)/a^{2}(\xi)\rm d\xi}.$ (19)
Consequently, the formulation of the forward diffusion process is equivalent to the definition of diffusion kernels characterized by $a(s)$ and $\sigma(s)$.
Conventional denoising diffusion probabilistic models (DDPM) [16] can be seen as discretizations of the variance preserving SDE with $a(s)=\sqrt{1-\beta(s)}$ and $\sigma(s)=\sqrt{\beta(s)/{(1-\beta(s))}}$, as $a^{2}(s)+a^{2}(s)\sigma^{2}(s)=1$. However, in order to achieve more efficient generation in GODM, we adopt EDM [15], which belongs to variance exploding SDE. Variance exploding SDE set $a(s)=1$, which implies that noise is directly added to the data instead of being blended through weighting. In this case, the variance of the noise (the noise level) is exclusively determined by $\sigma(s)$. In EDM, a linear noise level $\sigma(s)=s$ is applied. Therefore, the diffusion kernel can be written as:
@@
Reverse Denoising Process. As derived in [23], the reverse denoising process is formulated as the reverse SDE of Eq. 15:
@@
where $\nabla_{\mathbf{Z}}\log p_{s}(\mathbf{Z})$ is the score function of $\mathbf{Z}$. In EDM, given $a(s)=1$ and Eq. 18, it follows that,
$\bm{f}(\mathbf{Z},s)=f(s)\mathbf{Z}=\bm{0}.$ (23)
For the $g(s)$, derived from Eq. 19, we obtain
$g(s)=\sqrt{2\sigma(s)\dot{\sigma}(s)},$ (24)
@@
III-D Training
Graph Clustering. When the graph scales to millions of nodes, full-batch training becomes impractical, even with negative sampling. In addition, traditional neighbor sampling methods are well-suited in our case, as we need to reconstruct both the node feature and graph structure (i.e., edges). Specifically, node sampling cannot gather complete information for edge prediction, while edge sampling favors the nodes with a high degree. To address these challenges, we resort to graph clustering inspired by [25]. We first apply graph clustering algorithms (e.g., Metis [26]) to divide the large graph into small partitions. Then, we treat each partition as a minibatch and train GODM on each partition instead of the full graph.
Variational Encoder and Graph Generator. By integrating Variational Encoder and Graph Generator, we are able to train both of them in a variational autoencoder (VAE) fashion. We reconstruct the node feature with mean squared error:
$\ell_{\mathbf{X}}=\frac{1}{|\mathcal{V}|}\sum_{v_{i}\in\mathcal{V}}\|\mathbf{x}_{i}-\hat{\mathbf{x}}_{i}\|_{2}^{2},$ (26)
where $\|\cdot\|_{2}$ denotes the L2 norm. Then, we predict the edges in the training edge set $\bar{\mathcal{E}}$ with binary cross entropy loss:
$\ell_{\mathcal{E}}=-\frac{1}{|\bar{\mathcal{E}}|}\sum_{\bar{e}_{ij}\in\bar{\mathcal{E}}}(\bar{e}_{ij}\log(\hat{e}_{ij})+(1-\bar{e}_{ij}\log(1-\hat{e}_{ij}))).$ (27)
Similarly, we reconstruct the timestamp with mean squared error and edge type with cross-entropy loss:
$\ell_{\mathbf{t}}=\frac{1}{|\mathcal{E}|}\sum_{e_{ij}\in\mathcal{E}}(t_{ij}-\hat{t}_{ij})^{2},$ (28)
$\ell_{\mathbf{p}}=-\frac{1}{|\mathcal{E}|}\sum_{e_{ij}\in\mathcal{E}}\log(\hat{p}_{ij,p_{ij}}),$ (29)
We can obtain the total reconstruction loss by the weighted sum of all reconstruction loss:
$\ell_{\rm recon}=\omega_{\mathbf{X}}\ell_{\mathbf{X}}+\omega_{\mathcal{E}}\ell_{\mathcal{E}}+\omega_{\mathbf{t}}\ell_{\mathbf{t}}+\omega_{\mathbf{p}}\ell_{\mathbf{p}},$ (30)
where $\omega_{\mathbf{X}}$, $\omega_{\mathcal{E}}$, $\omega_{\mathbf{t}}$, and $\omega_{\mathbf{p}}$ are the hyperparameter weights balancing the importance between each term of reconstruction loss. Finally, we use the ELBO loss to train Variational Encoder and Graph Generator:
$\mathcal{L}_{\text{VAE}}=\ell_{\rm recon}+\beta\ell_{\rm kl}.$ (31)
where the $\ell_{\rm kl}$ is the Kullback-Leibler (KL) divergence loss between the latent space embedding and prior distribution (e.g., Gaussian distribution). However, as we have an additional Latent Diffusion Model, we loosen this regularization on the latent embedding. Consequently, we use a small hyperparameter weight $\beta<1$ to encourage the model to minimize reconstruction error while ensuring that the resultant embedding shape remains within the desired shape. The training of the VAE, consisting of Variational Encoder and Graph Generator, is presented in Algorithm 1.
@@ III-D Training
4:     Reparameterization of $\mathbf{Z}$ via Eq. 8
5:     Generate $\hat{\mathbf{X}}$, $\hat{\mathcal{E}}$, $\hat{\mathbf{t}}$, $\hat{\mathbf{p}}$ via Eq. 10, 12, 14, 13
6:     Calculate loss via Eq. 31
@@
Latent Diffusion Model. After the training of Variational Encoder and Graph Generator, we train Latent Diffusion Model via denoising score matching.
To solve the SDE in Eq. 25, we need to obtain the score function $\nabla_{\mathbf{Z}}\log p_{s}(\mathbf{Z})$, which is as intractable as $p_{s}(\mathbf{Z})=p(\mathbf{Z}^{s})$. However, the conditional distribution $p(\mathbf{Z}^{s}|\mathbf{Z}^{0})$ is tractable. From Eq. 17, we obtain its analytical solution:
Thus, we train a multi-layer perceptron (MLP) as the denoising function denoted by $\bm{\epsilon}_{\theta}$ for denoising score matching:
where ${\mathbf{Z}}^{s}_{C}=\mathbf{Z}^{s}+\mathbf{y}\mathbf{W}^{C},\mathbf{W}^{C}\in\mathbb{R}^{1\times d^{L}}$, encoding class label for conditional generation. The entire training process of Latent Diffusion Model is summarized in Algorithm 2.
@@
III-E Inference
The well-trained GODM is utilized in the inference process, augmenting real organic data by generated data. We first sampling $\hat{\mathbf{Z}}^{{S}}$ from Gaussian distribution. Then, we iteratively denoise $\hat{\mathbf{Z}}^{{s_{i}}}$ to obtain $\hat{\mathbf{Z}}^{{s_{i-1}}}$ conditioned on $\hat{\mathbf{y}}=\bm{1}$ via denoising function $\bm{\epsilon}_{\theta}$. The estimated $\hat{\mathbf{Z}}^{{0}}$ is fed into Graph Generator along with $\hat{\mathbf{y}}$ to generate synthetic graph $\hat{\mathcal{G}}$. Finally, the augmented graph $\mathcal{G}_{\text{sub}}$, which integrates the synthetic graph $\hat{\mathcal{G}}$ and the real organic graph $\mathcal{G}$, is used for training the downstream outlier detector $D$. Algorithm 3 presents the whole procedure of inference.
@@ III-F Complexity Analysis
In this section, we meticulously analyze and elucidate the time efficiency and memory scalability of GODM in training and inference for both VAE (Variational Encoder and Graph Generator) and Latent Diffusion Model, providing insights for handling large graphs which are common in real-world applications of graph outlier detection.
III-F1 Time Efficiency
@@
In the training of VAE, the major bottleneck comes from the graph structure generation in Graph Generator. This generation originally requires link prediction for every pair of nodes in the graphs, which is $O(n^{2})$. By graph clustering, we divide the large graph into relatively small partitions for minibatch training. We denote the average partition size (i.e., batch size) by $b$, so the number of partitions is $\frac{n}{b}$. Then, the complexity is reduced to $O(\frac{n}{b}b^{2})=O(nb)$, as only node pairs within each partition are predicted. We apply the negative sampling to further reduce the complexity to $O(|\mathcal{E}|)$, as we only sample a negative edge for each positive edge. For Latent Diffusion Model, we only sample a fixed number of diffusion steps to train the denoising function. As a result, assuming the number of latent dimensions is a constant, the complexity is $O(n)$. Thus, the total training complexity is $O(n+|\mathcal{E}|)$. For the inference, we first take $S$ steps sampling in Latent Diffusion Model, which is $O(nS)$, and predict edges between node pairs in each partition, which is $O(nb)$. Therefore, the total inference complexity is $O(nS+nb)$.
III-F2 Memory Scalability
@@
The vanilla VAE requires memory consumption of $O(n^{2})$, which is infeasible for large graphs. To reduce memory consumption, we adopt minibatch training on graph partitions and negative sampling for link prediction. The complexity of VAE training can be obtained by the number of edges divided by the number of partitions $O(|\mathcal{E}|/\frac{n}{b})=O(\frac{|\mathcal{E}|b}{n})$. Latent Diffusion Model only requires a memory of $O(b)$. In the inference process, the VAE performs link prediction for all node pairs in a batch, leading to a memory complexity of $O(b^{2})$, and Latent Diffusion Model only demands $O(b)$ for each batch. Therefore, the total inference process needs $O(b^{2})$ of memory.
IV Experiments
In this section, we systematically delineate the experiments to evaluate GODM. We first introduce experimental setups in Section IV-A. Then, we aim to answer: RQ1 (Section IV-B): How effective is GODM on improving graph outlier detection performance? RQ2 (Section IV-C): How is the quality of the synthetic data generated by GODM? RQ3 (Section IV-D): How efficient is GODM in terms of time and memory? The code implementation of our experiments is publicly available at https://github.com/kayzliu/godm.
IV-A Experimental Setups
To benchmark GODM and contemporary graph outlier detection algorithms from various aspects, we conduct experiments in a unified and comprehensive environment, which is in part adapted from GADBench [6].
IV-A1 Datasets
TABLE I: Statistics of datasets
@@ IV-A1 Datasets
(*: with multiple edge types and temporal infomation).
Dataset #Nodes #Edges #Feat. Outlier
@@
We use five diverse datasets from GADBench. Weibo, Elliptic, Tolokers, and Questions are homogeneous static graphs with only one type of edge. DGraph is a heterogeneous temporal graph with different types of edges. The graphs vary in scale, ranging from thousands of nodes to millions of nodes. Table I provides the statistics of the datasets. In the table, #Nodes stands for the number of nodes, and #Edges stands for the number of edges. #Feat. denotes the raw feature dimension, i.e., the number of node attributes. The Outlier column represents the outlier ratio in the label, indicating the extent of class imbalance. The detailed descriptions for each dataset are as follows:
@@
Tolokers [28]: The dataset is obtained from the Toloka crowdsourcing platform. It is composed of nodes corresponding to individual workers who have engaged in at least one out of thirteen specified projects. Edges are established between pairs of workers who have concurrently contributed to an identical task. The primary objective is to predict the likelihood of a worker having received a ban in any one of the projects. The features attributed to each node are constructed utilizing the worker’s personal profile information in conjunction with their task-related performance metrics.
Questions [28]: The dataset is derived from Yandex Q, a question-answering platform. The nodes represent individual users. An edge is established between two nodes to denote the scenario wherein one user has responded to another’s inquiry within the temporal bounds of one year, stretching from September 2021 to August 2022. This dataset is specifically curated to encapsulate the engagement of users who exhibit interest in the medical topic. The target for this dataset is to predict the likelihood of users’ continued activity on the platform by the conclusion of the observed period. The average FastText embeddings of the lexical content present in the users’ descriptions are used for node features.
@@
DGraph [7]: The DGraph dataset is a large-scale graph with different edge types and temporal information supplied by Finvolution Group. It includes 3 million nodes, 4 million dynamic edges, and 1 million labeled nodes. The nodes represent user accounts within a financial organization offering personal loan services, while an edge between two nodes indicates that one account has designated the other as an emergency contact. Nodes classified as fraud correspond to users displaying delinquent financial conduct. For those accounts with borrowing records, outliers are identified as accounts with a history of overdue payments, while inliers are those without such a history. Additionally, the dataset includes 17 node features derived from user profile information.
IV-A2 Baselines
In our experiments, we evaluate GODM against different types of baseline methods. We have three types of baseline methods, including general graph neural networks, graph outlier detectors, and data augmentation for graph outlier detection. For general graph neural networks, we follow GADBench and select six GNNs:
@@ IV-A2 Baselines
In addition to general GNNs, we also include eight GNNs specifically designed for graph outlier detection:
@@
DCI (Deep Cluster Infomax) [34]: DCI is a self-supervised learning strategy that separates the learning of node representations from outlier detection. It addresses discrepancies between node behavioral patterns and their label semantics by clustering the entire graph, thus capturing intrinsic graph properties in focused feature spaces.
@@
BWGNN (Beta Wavelet Graph Neural Network) [39]: BWGNN addresses the ”right-shift” phenomenon on outliers. It uses the Beta kernel to address higher frequency anomalies through spatially/spectrally localized band-pass filters.
@@
For heterogeneous graphs, despite the rich literature, limited heterogeneous GNNs can cope with temporal information. As one of the representatives, we chose HGT for our baseline:
HGT (Heterogeneous Graph Transformer) [21]: HGT is engineered to address the challenges of modeling heterogeneous graphs. It introduces node- and edge-type dependent parameters that are crucial for defining heterogeneous attention mechanisms for each edge.
Because the data augmentation for graph outlier detection is a relatively new research topic, there are limited baselines. We include one baseline for homogeneous graphs and introduce a variant of GODM that can work on both homogeneous and heterogeneous datasets, GOVAE:
DAGAD (Data Augmentation for Graph Anomaly Detection) [41]: DAGAD incorporates three modules to augment the graph data, including an information fusion module for representation learning, a data augmentation module to enrich the training set with synthetic samples, and an imbalance-tailored learning module to distinguish between minority anomalous class and majority normal class.
GOVAE: In this variant, we drop Latent Diffusion Model and form a variational autoencoder with Variational Encoder and Graph Generator. The inference is a direct one-step estimation from prior Gaussian distribution.
Along with GODM, we compare these three data augmentation methods for graph outlier detection in Table II, where the flexibility represents whether the data augmentation is model agnostic. Heterogeneous and time indicate whether the method supports these two types of information. The diffusion column indicates whether the method adopts diffusion models.
TABLE II: Comparison of data augmentation for graph outlier detection.

### v3

@@ B.1 Forward Diffusion Process
We construct a forward diffusion process $\{\mathbf{Z}(s)\}_{s=0}^{S}$, where $s\in[0,S]$ is a continuous diffusion step. In the diffusion process, $\mathbf{Z}(0)=\mathbf{Z}$ is the embedding from Variational Encoder, while $\mathbf{Z}(S)\sim\mathcal{N}(\bm{0},\bm{I})$ is sampled from the prior distribution. According to [52], the forward diffusion process can be written in stochastic differential equation (SDE) as:
@@
where $\bm{\omega}_{t}$ is the standard Wiener process (i.e., Brownian motion). $\bm{f}(\cdot,s)$ and $g(\cdot)$ are the drift coefficient and the diffusion coefficient, respectively. The selection of $\bm{f}(\cdot,s)$ and $g(\cdot)$ vary between different diffusion models. $f(\cdot)$ is usually of the form $\bm{f}(\mathbf{Z},s)=f(s)\mathbf{Z}$. Thus, Equation 10 can be written as:
@@
Let $\mathbf{Z}$ be a function of diffusion step $s$, i.e., $\mathbf{Z}^{s}=\mathbf{Z}(s)$. The diffusion kernel of Equation 11 can be represented in the conditional distribution of $\mathbf{Z}^{s}$ given $\mathbf{Z}^{0}$:
@@
$a(s)=\exp{\int_{0}^{s}f(\xi)\rm d\xi},\quad\sigma(s)=\sqrt{\int_{0}^{s}g^{2}(\xi)/a^{2}(\xi)\rm d\xi}.$ (13)
Consequently, the formulation of the forward diffusion process is equivalent to the definition of diffusion kernels characterized by both $a(s)$ and $\sigma(s)$.
Denoising diffusion probabilistic models (DDPM) [13] can be seen as discretizations of the variance preserving SDE with $a(s)=\sqrt{1-\beta(s)}$ and $\sigma(s)=\sqrt{\beta(s)/{(1-\beta(s))}}$, as $a^{2}(s)+a^{2}(s)\sigma^{2}(s)=1$. However, in order to achieve more efficient generation in GODM, we adopt EDM [12], which belongs to variance exploding SDE. Variance exploding SDE set $a(s)=1$, which implies that noise is directly added to the data instead of being blended through weighting. In this case, the variance of the noise (the noise level) is exclusively determined by $\sigma(s)$. In EDM, a linear noise level $\sigma(s)=s$ is applied. Therefore, the diffusion kernel can be written as:
@@
B.2 Reverse Denoising Process
As derived in [52], the reverse denoising process is formulated as the reverse SDE of Equation 10:
@@ B.2 Reverse Denoising Process
where $\nabla_{\mathbf{Z}}\log p_{s}(\mathbf{Z})$ is the score function of $\mathbf{Z}$. In EDM, given $a(s)=1$ and Equation 13,
$\bm{f}(\mathbf{Z},s)=f(s)\mathbf{Z}=\bm{0},\quad g(s)=\sqrt{2\sigma(s)\dot{\sigma}(s)},$ (17)
@@
B.3 Training of Latent Diffusion Models
To solve the SDE in Equation 18, we need to obtain the score function $\nabla_{\mathbf{Z}}\log p_{s}(\mathbf{Z})$, which is as intractable as $p_{s}(\mathbf{Z})=p(\mathbf{Z}^{s})$. However, the conditional distribution $p(\mathbf{Z}^{s}|\mathbf{Z}^{0})$ is tractable. From Equation 12, we obtain its analytical solution:
Appendix C Algorithms
In this section, we provide algorithms used in GODM. Algorithm 1 gives the training process of VAE. Algorithm 2 shows the training process of Latent DIffusion Model. Algorithm 3 presents the inference process.
@@ Appendix C Algorithms
4:     Reparameterization of $\mathbf{Z}$ via Equation 3
5:     Generate $\hat{\mathbf{X}}$, $\hat{\mathcal{E}}$, $\hat{\mathbf{t}}$, $\hat{\mathbf{p}}$ via Equation 4, 5, 7, 6
6:     Calculate loss via Equation 8
@@ Appendix D Complexity Analysis
In this section, we meticulously analyze and elucidate the time efficiency and memory scalability of GODM in training and inference for both VAE (Variational Encoder and Graph Generator) and Latent Diffusion Model.
D.1 Time Efficiency
@@
In the training of VAE, the major bottleneck comes from the graph structure generation in Graph Generator. This generation originally requires link prediction for every pair of nodes in the graphs, which is $O(n^{2})$. By graph clustering, we divide the large graph into relatively small partitions for mini-batch training. We denote the average partition size (i.e., batch size) by $b$, so the number of partitions is $\frac{n}{b}$. Then, the complexity is reduced to $O(\frac{n}{b}b^{2})=O(nb)$, as only node pairs within each partition are predicted. We apply the negative sampling to further reduce the complexity to $O(|\mathcal{E}|)$, as we only sample a negative edge for each positive edge. For Latent Diffusion Model, we only sample a fixed number of diffusion steps to train the denoising function. As a result, assuming the number of latent dimensions is a constant, the complexity is $O(n)$. Thus, the total training complexity is $O(n+|\mathcal{E}|)$. For the inference, we first take $S$ steps sampling in Latent Diffusion Model, which is $O(nS)$, and predict edges between node pairs in each partition, which is $O(nb)$. Therefore, the total inference complexity is $O(nS+nb)$.
D.2 Memory Scalability
@@
The vanilla VAE requires memory consumption of $O(n^{2})$, which is infeasible for large graphs. To reduce memory consumption, we adopt mini-batch training on graph partitions and negative sampling for link prediction. The complexity of VAE training can be obtained by the number of edges divided by the number of partitions $O(|\mathcal{E}|/\frac{n}{b})=O(\frac{|\mathcal{E}|b}{n})$. Latent Diffusion Model only requires a memory of $O(b)$. In the inference process, the VAE performs link prediction for all node pairs in a batch, leading to a memory complexity of $O(b^{2})$, and Latent Diffusion Model only demands $O(b)$ for each batch. Therefore, the total inference process needs $O(b^{2})$ of memory.
Appendix E Implementation Details
We modified GADBench [3] to benchmark graph outlier detection performance.
Environment. The key libraries and their versions used in experiments are as follows: Python 3.9, CUDA 11.8, PyTorch 2.0.1 [53], PyG 2.4.0 [54], DGL 1.1.2 [55], and PyGOD 1.0.0 [1].
Hyperparameters. GODM is implemented with default hyperparameters. For Variational Encoder, we use one layer GraphSAGE for $\text{GNN}_{\text{shared}}$, $\text{GNN}_{\bm{\mu}}$, and $\text{GNN}_{\bm{\sigma}}$, respectively. We set the hidden dimension to the largest power of 2 that is no greater than the feature dimension of the dataset divided by 2. To avoid extensive hyperparameter search, we set the weights of reconstruction loss $\omega_{\mathbf{X}}$, $\omega_{\mathcal{E}}$, $\omega_{\mathbf{t}}$, and $\omega_{\mathbf{p}}$ to roughly balance their absolute value only, which are 1, 0.5, 1, and 0.3, respectively. The weight for KL-divergence $\beta$ is 0.001. For the diffusion model, we use the dimension twice as the hidden dimension in VAE. We use five layers of MLP as the denoising function following the detailed design in [51]. For training, we adopt the Adam optimizer with a learning rate of 0.001 without weight decay to train the VAE and the diffusion model for 100 epochs, respectively, and apply early stopping with a patience of 50. The negative sampling ratio is 1, and the approximate graph partition size is 2048. For inference, we use 50 steps of diffusion and generate the same amount of synthetic outliers as the number of real organic outliers in the training set. For all the hyperparameters in graph outlier detection, we apply the default setting in GADBench. We report the graph outlier detection performance of GODM with the optimal downstream graph outlier detector.
Appendix F Description of Datasets
@@
Table 4: Statistics of datasets (*: with multiple edge types and temporal infomation).
Dataset #Nodes #Edges #Features Outlier
@@
This section describes the datasets used in the experiments. Table 4 provides the statistics of the datasets. In the table, #Nodes stands for the number of nodes, and #Edges stands for the number of edges. #Feature denotes the raw feature dimension, i.e., the number of node attributes. The Outlier column represents the outlier ratio in the label, indicating the extent of class imbalance. The detailed descriptions for each dataset used in the experiments are as follows:
@@
Tolokers [57]: The dataset is obtained from the Toloka crowdsourcing platform. It is composed of nodes corresponding to individual workers who have engaged in at least one out of thirteen specified projects. Edges are established between pairs of workers who have concurrently contributed to an identical task. The primary objective is to predict the likelihood of a worker having received a ban in any one of the projects. The features attributed to each node are constructed utilizing the worker’s personal profile information in conjunction with their task-related metrics.
Questions [57]: The dataset is derived from Yandex Q, a question-answering platform. The nodes represent users, and an edge is established between two nodes to denote the scenario wherein one user has responded to another’s inquiry within the temporal bounds of one year, stretching from September 2021 to August 2022. This dataset is specifically curated to encapsulate the engagement of users who exhibit interest in the medical topic. The target for this dataset is to predict the likelihood of users’ continued activity on the platform by the conclusion of the observed period. The average FastText embeddings of the lexical content present in the users’ descriptions are used for node features.
@@
DGraph [4]: DGraph is a large-scale graph with different edge types and temporal information supplied by Finvolution Group. It includes around 3 million nodes, 4 million dynamic edges, and 1 million node labels. The nodes represent user accounts within a financial organization offering personal loan services, while an edge between two nodes indicates that one account has designated the other as an emergency contact. Nodes classified as fraud correspond to users displaying delinquent financial conduct. For those accounts with borrowing records, outliers are identified as accounts with a history of overdue payments, while inliers are those without such a history. Additionally, the dataset includes 17 node features derived from user profile information.
Appendix G Description of Baselines
This section provides detailed descriptions of the baselines in the experiments.
G.1 General Graph Neural Network
@@
HGT (Heterogeneous Graph Transformer) [18]: HGT is designed to address the challenges of modeling heterogeneous graphs. It introduces node- and edge-type dependent parameters for heterogeneous attention mechanisms for each type of edge.
G.2 Graph Outlier Detector
@@
DCI (Deep Cluster Infomax) [28]: DCI is a self-supervised learning strategy that separates the learning of node representations from outlier detection. It addresses discrepancies between node behavioral patterns and their label semantics by clustering, thus capturing intrinsic graph properties in focused feature spaces.
@@
BWGNN (Beta Wavelet Graph Neural Network) [33]: BWGNN addresses the "right-shift" phenomenon on outliers. It uses the Beta kernel to address higher frequency anomalies through spatially/spectrally localized band-pass filters.
@@
G.3 Data Augmentation
DAGAD (Data Augmentation for Graph Anomaly Detection) [35]: DAGAD incorporates three modules to augment the graph, including an information fusion module for representation learning, a data augmentation module to enrich the training set with synthetic samples, and an imbalance-tailored learning module to distinguish between minority anomalous class and majority normal class.
GOVAE: In this variant, we drop Latent Diffusion Model and form a variational autoencoder with Variational Encoder and Graph Generator. The inference process is a direct one-step estimation from prior Gaussian distribution.
Along with GODM, we compare three data augmentation methods in Table 5, where the flexibility represents whether the data augmentation is model agnostic. Heterogeneous and time indicate whether the method supports these two types of information. The diffusion column indicates whether the method adopts diffusion models.
Table 5: Comparison of data augmentation methods.